OpenStudy (solomonzelman):

This is physics, really. (sorry for posting here) Do I have a brain injury?

OpenStudy (sbuck98):

Explain your condition, I'll help.

OpenStudy (solomonzelman):

This relates to projectile motion range. REMINDER: Projectile path is symmetric about the vertex, and vertex is the maximum. So, I know that \(\color{#000000 }{ \displaystyle v_{y}(t)=v_{iy} -gt}\) (velocity in the y-direction) To find an expression for vertical displacement/position at time t, I integrate both sides, \(\color{#000000 }{ \displaystyle \int v_{y}(t)~dt=\int~v_{iy} -gt~dt}\) \(\color{#000000 }{ \displaystyle y(t)=v_{iy}t -\frac{1}{2}gt^2+y_0}\) (In this context, +C is the initial vertical position from where the particle starts to travel) ~~~~~~~~~~~~~~~~~~~~~~~~~~~ Then, I know that the time to reach the top (recall that vertical velocity at the top is zero), \(\color{#000000 }{ \displaystyle 0=v_{iy} -gt}\) \(\color{#000000 }{ \displaystyle gt=v_{iy} }\) \(\color{#000000 }{ \displaystyle t=v_{iy}/g }\) (time to get to the top) Or, in terms of the initial angle and initial speed, \(\color{#000000 }{ \displaystyle y(t)=v_{i}\sin(\theta)t -\frac{1}{2}gt^2+y_0}\) (the vertical position as a function of time) \(\color{#000000 }{ \displaystyle t_{\rm ~top}=v_{i}\sin(\theta)/g }\) (time to reach the top) So the maximum y, (the range) is, \(\color{#000000 }{ \displaystyle {\rm Range}=y(v_{i}\sin(\theta))}\) And for that I get, \(\color{#000000 }{ \displaystyle {\rm Range}=v_{i}\sin(\theta)\left(\frac{v_{i}\sin(\theta)}{g}\right) -\frac{1}{2}g\left(\frac{v_{i}\sin(\theta)}{g}\right)^2+y_0}\) \(\color{#000000 }{ \displaystyle {\rm Range}=\frac{v_{i}^{~2}\sin^2(\theta)}{g} -\frac{1}{2}\frac{v_{i}^{~2}\sin^2(\theta)}{g}+y_0}\) \(\color{#000000 }{ \displaystyle {\rm Range}=\frac{1}{2}\frac{v_{i}^{~2}\sin^2(\theta)}{g}+y_0}\) \(\color{#000000 }{ \displaystyle {\rm Range}=\frac{v_{i}^{~2}\sin^2(\theta)}{2g}+y_0}\)

OpenStudy (solomonzelman):

But, the formula says \(\color{#000000 }{ \displaystyle {\rm Range}=\frac{v_{i}^{~2}\sin (2\theta)}{g}}\) (well the initial vertical positon \(y_0\) is not that important)

OpenStudy (thomas5267):

Shouldn't \(t_{top}=\dfrac{v_i\sin(\theta)}{2g}\)?

OpenStudy (solomonzelman):

No, there is \(2\theta\) there (in the formula).

OpenStudy (solomonzelman):

and the initial speed is squared.

OpenStudy (solomonzelman):

Oh, that formula is for horizontal range, and mine is vertical

OpenStudy (solomonzelman):

I just need to verify if my result is indeed the correct maximum height...

OpenStudy (solomonzelman):

Oh, nvm I got it correctly... I seee the formula online ;)

OpenStudy (solomonzelman):

I guess there isn't a question, sorry for this waste. Should have webbed carefully before

OpenStudy (tkhunny):

In response to your initial question, it appears the answer is "no".

OpenStudy (solomonzelman):

Oh, lol, tnx:)

OpenStudy (solomonzelman):

I'll lear a lesson. Search Search and Search again, then ask. (I just have that connotation that Range is something vertical)

OpenStudy (thomas5267):

I just checked my notes. Yes the answer is correct and the range is for the horizontal range not the maximum vertical height.

OpenStudy (solomonzelman):

Yes:)

OpenStudy (solomonzelman):

Time to close:)

OpenStudy (solomonzelman):

tnx

OpenStudy (solomonzelman):

I will just finish the horizontal tange derivation for myself to leave with a good mind.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle v_x(t)=v_{i}\cos(\theta) }\) \(\color{#000000 }{ \displaystyle \int v_x(t)~dt=\int v_{i}\cos(\theta)~dt }\) \(\color{#000000 }{ \displaystyle x(t)=v_{i}\cos(\theta)t}\) (wil omit the constant, or the initial position in the x, since we always choose \(x_0=0\), by projectile motion. at least, I don't recal when this wasn't the case) time to reach the top is (previously derived) \(\color{#000000 }{ \displaystyle t_{\rm top}=\frac{v_{i}\sin(\theta)}{g}}\) projectile path is symmetric about vertex/top, so \(\color{#000000 }{ \displaystyle t_{\rm ~flight}=t_{\rm ~top}=\frac{2v_{i}\sin(\theta)}{g}}\) Horizontal range, \(\color{#000000 }{ \displaystyle {\rm R_x}=x(\frac{2v_{i}\sin(\theta)t}{g})=v_{i}\cos(\theta)\frac{2v_{i}\sin(\theta)t}{g}}\) \(\color{#000000 }{ \displaystyle {\rm R_x}=\frac{2v_{i}^{~2}\sin(\theta)\cos(\theta)}{g}=\frac{v_{i}^{~2}\sin(2\theta)}{g}}\)

OpenStudy (solomonzelman):

boom boom :)